if 300. mL of 1.50 M Mg(nNO3)2 produces 400. mL of AgNO3, what is the concentration in M?
Answer #1

Applying molarity equation M_{1}V_{1} = M_{2}V_{2}

M1 = 1.50M , V1 = 300 ml

M2= ? ,V2 =400 ml

1.50\times 300 = M_{2}\times400

M_{2}= 450/400 = 1.125 M